# Around the ‘‘Fundamental Theorem of Algebra’’

### v1

@@ Abstract.
This research was carried out at the Higher School of Contemporary Mathematics of Moscow Institute of Physics and Technology, with the support of the Ministry of Science and Higher Education of the Russian Federation, project no. SMG - 2024 -0048.
@@ 1. Kac Theorem
A version of the Fundamental Theorem of Algebra (FTA) concerning the real roots of a real polynomial, based on the concept of the probability that a root is real (sometimes referred to as the Kac theorem), was established in 1938 by J. Littlewood and A. Offord; see [1, 2, 3, 4]. The origin of the Kac theorem lies in the following simple observation: for any integer $m$, the probability that a root of a real polynomial of degree $m$ is real is nonzero. For instance, if $m=1$, this probability equals 11.
@@ 2.1. Formulation of the FTA
(1) A Laurent polynomial $P$ is called real if its values on the circle $S$ of radius 11 centered at the origin are real.
@@
(3) $\kappa(S)$ belongs to a sphere of radius 11 centered at zero in ${\rm Trig}(\Lambda)$
@@ 3.1. Exponential sums of one variable
Suppose, to the contrary, that the constant $C(A)$ does not exist. Then there exists a sequence of points $w_{1},w_{2},\ldots$ in $\mathbb{C}$} such that the number of zeros of the function $f(w_{k}+z)$ in the set AA tends to infinity as $k\to\infty$. We can choose this sequence so that the corresponding sequence $\rho(w_{k})(f)$ converges in $\mathbb{P}$} to a limit $f_{\infty}$. However, this would imply that the exponential sum $f_{\infty}$ has infinitely many zeros in AA, which is impossible. ∎
@@ 3.2. Exponential sums of several variables
    $A(\Gamma)$ is the exterior angle of $\Gamma$, i.e., the angle of the dual cone $K_{\Gamma}$ corresponding to the face $\Gamma$ (the full $n$-dimensional angle is normalized to 11). Recall that $K_{\Gamma}$ consists of those points $w\in{\mathbb{C}}^{n}$ for which the linear functional $f(z)=\Re\,z(w)$, considered as a function on $\Delta$, attains its maximum simultaneously at all points of the face $\Gamma$.

### v2

@@ 1. Kac Theorem
A version of the Fundamental Theorem of Algebra (FTA) concerning the real roots of a real polynomial, based on the concept of the probability that a root is real (sometimes referred to as the Kac theorem), was established in 1938 by J. Littlewood and A. Offord; see [1, 2, 3, 4]. The origin of the Kac theorem lies in the following simple observation: for any integer $m$, the probability that a root of a real polynomial of degree $m$ is real is nonzero. For instance, if $m=1$, this probability equals $1$.
@@ 2.1. Formulation of the FTA
(1) A Laurent polynomial $P$ is called real if its values on the circle $S$ of radius $1$ centered at the origin are real.
@@ 2.2. Proof of Theorem 3
(3) $\kappa(S)$ belongs to a sphere of radius $1$ centered at zero in ${\rm Trig}(\Lambda)$
@@ 3.1. Exponential sums of one variable
Suppose, to the contrary, that the constant $C(A)$ does not exist. Then there exists a sequence of points $w_{1},w_{2},\ldots$ in $\mathbb{C}$} such that the number of zeros of the function $f(w_{k}+z)$ in the set $A$ tends to infinity as $k\to\infty$. We can choose this sequence so that the corresponding sequence $\rho(w_{k})(f)$ converges in $\mathbb{P}$} to a limit $f_{\infty}$. However, this would imply that the exponential sum $f_{\infty}$ has infinitely many zeros in $A$, which is impossible. ∎
@@ 3.2. Exponential sums of several variables
    $A(\Gamma)$ is the exterior angle of $\Gamma$, i.e., the angle of the dual cone $K_{\Gamma}$ corresponding to the face $\Gamma$ (the full $n$-dimensional angle is normalized to $1$). Recall that $K_{\Gamma}$ consists of those points $w\in{\mathbb{C}}^{n}$ for which the linear functional $f(z)=\Re\,z(w)$, considered as a function on $\Delta$, attains its maximum simultaneously at all points of the face $\Gamma$.